# 03F40 (primary), 03D20

### v2

@@ 7 Independence
If $s=b$, then $U(s)={\Omega}$ and $U(s)[\iota]=\iota$, so since $\iota=u_{B}(c)$ with $c<b$ in $u_{B}$-normal form, we may take $s^{\prime}=c$. Then, since $c<b$, $U(s^{\prime})=u_{B}(c)$, and
$U(c)=u_{B}(c)=\iota={\Omega}[\iota]=U(b)[\iota].$
Otherwise, write $s=_{b}b^{e}a+r$ and $U(s)=_{\Omega}{\Omega}^{\eta}{\alpha}+\gamma$. If $\gamma>0$, then $U(r)=\gamma$. We apply the induction hypothesis to find suitable $r^{\prime}<r$ and set $s^{\prime}={b}^{e}a+r^{\prime}$. Since $U(r^{\prime})<U(r)$ and ${\Omega}^{U(e)}u_{B}(a)+U(r)$ is in $\Omega$}-normal form, we easily deduce that ${\Omega}^{U(e)}u_{B}(a)+U(r^{\prime})$ is in $\Omega$}-normal form, from which we readily see that $s^{\prime}$ is in $U^{b}_{B}$-normal form.
If $\gamma=0$ but ${\alpha}>0$, suppose first that $\alpha$} is a limit. Then, $r=0$ and $U(a)=u_{B}(a)={\alpha}$. The assumption on digits yields $a^{\prime}\leq a$ in $u_{B}$-normal form such that ${\alpha}[\iota]=u_{B}(a)[\iota]=a^{\prime}$, and thus we may set $s^{\prime}=b^{e}a^{\prime}$ and see that
@@
If ${\alpha}=\delta+1>1$, then $({\Omega}^{\eta}{\alpha})[\iota]={\Omega}^{\eta}\delta+{\Omega}^{\eta}[\iota]$. We must then have that $u_{B}(a)=\delta+1$, which since $(\delta+1)[\iota]=\delta$ yields $a^{\prime}<a$ in $u_{B}$-normal form with $u_{B}(a^{\prime})=\delta$. We must also have that $U(b^{e})={\Omega}^{\eta}$, so the induction hypothesis yields suitable $r^{\prime}<b^{e}$ so that
@@
The $\Omega$}-decomposition condition is deduced from $U(r^{\prime})<U(r)$ as above. Finally, if ${\alpha}=1$, we see that $U(e)=\eta$, so we apply the induction hypothesis to find suitable $e^{\prime}$ with $u_{B}(e^{\prime})=u_{B}(e)[\iota]$. Let $a^{\prime}=1$ if $\eta$ is a limit, otherwise, $a^{\prime}=c$. We may then set $s^{\prime}=b^{e^{\prime}}a^{\prime}$, and once again, ${\Omega}^{U(e^{\prime})}u_{B}(a^{\prime})$ is in $\Omega$}-decomposition trivially.
@@
Consider transformations from a base hierarchy $B$ to $C=B_{+k}$, and let $i\leq\min(B\cup\{k+1\})$. If $n$ is in $u_{B}$-normal form, then there is $n^{\prime}<{\uparrow}n$ such that $u_{C}(n^{\prime})=u_{C}({\uparrow}n)[i]$ and $n^{\prime}$ is in $u_{C}$-normal form.
Proof 7.13.
Proceed by induction on $n$. Let $b={\rm Base}_{B}(n)$, $\xi=u_{C}({\uparrow}n)$, and $c={\rm Base}_{C}({\uparrow}n)$. If $n<\min B$, then we must have $\xi[i]=n-1$, so we can take $n^{\prime}=n-1$. If $n\in B$, then $\xi=\omega$ and we take $n^{\prime}=i=\omega[i]$.
Otherwise, let $\zeta=U_{B}(n)$. By Proposition 7.8, $\zeta=U_{C}({\uparrow}n)$, and $\xi=\psi(\zeta)$ is in $\psi$-normal form by Lemma 7.3.
If $\zeta$ has countable cofinality, $\xi[i]=\psi(\zeta[i])$. If $\zeta[i]<{\Omega}$, then Lemma 6.6 yields $\zeta[i]\in\{0,1\}$, hence $\xi[i]\in\{1,2\}$, and we may take $n^{\prime}=\xi[i]$, noting that $\min B\leq 2$, so $\min C\leq 3$. Otherwise, Lemma 7.10 yields $n^{\prime}<{\uparrow}n$ in $U^{c}_{C}$-normal form with $U^{c}_{C}(n^{\prime})=\zeta[i]$, and since $\zeta[i]\geq{\Omega}$, $n^{\prime}\geq c$, so ${\rm Base}_{C}(n^{\prime})=c$. Thus $n^{\prime}$ is in $U_{C}$-normal form, and since $\psi(\zeta[i])$ is in $\psi$-normal form by Lemma 6.51, Lemma 7.3 implies that $n^{\prime}$ is in $u_{C}$-normal form, as needed.
Otherwise, $\zeta$ has uncountable cofinality. Let $d_{j}=d_{j}(n)$ as in Definition 5.1. For $j\leq i$, we claim that there is $n_{j}<d_{j}$ such that $n_{j}$ is in $u_{C}$-normal form and $u_{C}(n_{j})=\xi[j]$.
For $j=0$, we have that $\xi[0]=0$ and we may take $n_{0}=0$. Otherwise, Lemma 7.6 yields
$U^{d_{j}}_{C}(\langle\begin{smallmatrix}d_{j}\\ b\end{smallmatrix}\rangle n)=U^{b}_{B}(n)=U_{B}(n)=\zeta.$
By induction on $j$, there is $n_{j}<d_{j}$ in $u_{C}$-normal form such that $u_{C}(n_{j})=\xi[j]$. By the main induction hypothesis, $U^{d_{j}}_{C}(\langle\begin{smallmatrix}d_{j}\\ b\end{smallmatrix}\rangle n)$ satisfies the assumptions of Lemma 7.10 regarding $d_{j}$-digits, so there is $n_{j+1}<\langle\begin{smallmatrix}d_{j}\\ b\end{smallmatrix}\rangle n=d_{j+1}$ such that $U^{d_{j}}_{C}(n_{j+1})=\zeta[\xi[j]]$ and $n_{j+1}$ is in $U^{d_{j}}_{C}$-normal form. If $\zeta[\xi[j]]<{\Omega}$, then $\zeta[\xi[j]]\in\{0,1\}$ by Lemma 6.6, and thus $n_{j+1}\in\{0,1\}$, accordingly, which is in $u_{C}$-normal form by definition. Otherwise, $\zeta[\xi[j]]\geq{\Omega}$, so $n_{j+1}\geq d_{j}$. Since $d_{j+1}=S_{C}(d_{j})$, we have that ${\rm Base}_{C}(n_{j+1})=d_{j}$, hence $n_{j+1}$ is in $U_{C}$-normal form. Moreover, since $\psi(\zeta[\xi[j]])$ is in $\psi$-normal form by Lemma 6.52, $n_{j+1}$ is in $u_{C}$-normal form by Lemma 7.3. We can thus take $n^{\prime}=n_{i}<d_{i}\leq d_{k+1}=\langle\begin{smallmatrix}d_{k}\\ b\end{smallmatrix}\rangle n={\uparrow}n$.
@@
To see that $n_{i}\leq m_{i}$ for all $i$, use Lemma 3.11 and induction to see that
$n_{i+1}\leq{\uparrow}_{i}n_{i}-1\leq{\uparrow}_{i}m_{i}-1=m_{i+1}.$
This implies that $m_{i}>0$ whenever $i<F(k)$. Thus $F$ is indeed bounded by $G$, so $G$ cannot be provably total in $\sf KP$ by Theorem 6.9. We conclude that Theorem 2.16 is also not provable in $\sf KP$.
8 Finite Base Hierarchies
The proof of Theorem 2.17 uses a dynamical hierarchy whose base hierarchies (aside from the first two) are infinite. One may wonder if the use of infinite base hierarchies is essential. The answer is no; here we will briefly sketch how our constructions should be modified to show this. Below, say that a finite dynamical hierarchy is a dynamical hierarchy $\mathcal{B}$ such that $\mathcal{B}_{i}$ is finite for all $i$.
Theorem 2.17

### v4

@@ 1. Case 1
  2. Case 2
    ($s\geq b$). Write $s=_{b}b^{e}a+r$ and $U(s)=_{\Omega}{\Omega}^{\eta}{\alpha}+\gamma$.
    1. Case 2.1
      ($\gamma>0$). Then, $U(r)=\gamma$. We apply the induction hypothesis to find suitable $r^{\prime}$ with either $r^{\prime}=r-1$ if $U(s)$ (hence $U(r)$) is a successor, and otherwise $r^{\prime}\leq r-\Delta$ if $U(r)$ is a limit, and set $s^{\prime}={b}^{e}a+r^{\prime}$, which respectively is either s1s-1 or at most $s-\Delta$. Since $U(r^{\prime})<U(r)$ and ${\Omega}^{U(e)}u_{B}(a)+U(r)$ is in $\Omega$}-normal form, we easily deduce that ${\Omega}^{U(e)}u_{B}(a)+U(r^{\prime})$ is in $\Omega$}-normal form, from which we readily see that $s^{\prime}$ is in $U^{b}_{B}$-normal form.
    2. Case 2.2
      ($\gamma=0$). In this case, $U(s)$ is a limit, so we only need to establish the second claim. Moreover, $U(a)=u_{B}(a)={\alpha}$. We consider cases according to $\alpha$}.
      1. Case 2.2.1
        ($\alpha$} is a limit). The assumption on digits yields $a^{\prime}\leq a-\Delta$ in $u_{B}$-normal form such that ${\alpha}[\iota]=u_{B}(a)[\iota]=a^{\prime}$, and thus we may set
        $s^{\prime}=b^{e}a^{\prime}\leq b^{e}a-b^{e}\leq b^{e}a-\min B\leq s-\Delta$
        (where the second inequality uses $e\geq 1$ since $s\geq b$) and see that
@@ 2. Case 2
      2. Case 2.2.2
        ($1<{\alpha}=\delta+1$). Then, $({\Omega}^{\eta}{\alpha})[\iota]={\Omega}^{\eta}\delta+{\Omega}^{\eta}[\iota]$. Let $a^{\prime}=a-1$ and $b^{\prime}={\rm Base}_{B}(a)\leq b$. Since $a^{\prime}+b^{\prime}<s+b$, the induction hypothesis applied to the first claim yields $U_{B}(a)=U^{b^{\prime}}_{B}(a)=U^{b^{\prime}}_{B}(a^{\prime})+1$. Since $U_{B}(a)$ is not a limit, by Lemma 7.2(1), $\min B\nmid a$, and thus $a\notin B$. It follows that $a^{\prime}\geq b^{\prime}$, so $b^{\prime}={\rm Base}_{B}(a^{\prime})$ and thus $U_{B}(a^{\prime})=U^{b^{\prime}}_{B}(a^{\prime})$, hence $u_{B}(a)=\delta$. We must also have that $U(b^{e})={\Omega}^{\eta}$, so the induction hypothesis yields suitable $r^{\prime}\leq b^{e}-\Delta$ so that $b^{e}a^{\prime}+r^{\prime}\leq s-\Delta$ and
@@
        The $\Omega$}-decomposition condition is deduced from $U(r^{\prime})<U(r)$ as above.
      3. Case 2.2.3
@@
        (${\alpha}=1$). We see that $U(e)=\eta$, so we apply the induction hypothesis to find suitable $e^{\prime}<e$ with $u_{B}(e^{\prime})=\eta[\iota]$. Let $a^{\prime}=1$ if $\eta$ is a limit, otherwise, $a^{\prime}=c$. We may then set $s^{\prime}=b^{e^{\prime}}a^{\prime}$, and once again, ${\Omega}^{U(e^{\prime})}u_{B}(a^{\prime})$ is in $\Omega$}-decomposition trivially. If $\eta$ is a limit, then $e^{\prime}\leq e-\Delta$, so
        $b^{e^{\prime}}\leq b^{e-\Delta}\leq b^{e}-\Delta,$
        where the second inequality is immediate from the fact that $x\mapsto b^{x}$ is strictly increasing. Otherwise,
        $b^{e^{\prime}}c\leq b^{e^{\prime}}(b-\Delta)\leq b^{e}-\Delta.\qed$
The general picture is that as we iteratively apply the upgrade from $\mathcal{D}_{i}$ to $\mathcal{D}_{i+1}$, we can compute more and more of the fundamental sequences. As a base case, we note that we can always compute them up to 0.
Lemma 7.10.
Let $B$ be any base hierarchy and $n>0$ be in $u_{B}$-normal form. Then, there is $n^{\prime}<n$ in $u_{B}$-normal form such that $u_{B}(n^{\prime})=u_{B}(n)[0]$. If, moreover, $u_{B}(n)$ is a limit, then we can take $n^{\prime}\leq n-\min B$.
Proof.
Proceed by induction on $n$. If $n<\min B$, then take $n^{\prime}=n-1$, and note that both are in $u_{B}$-normal form by definition. If $\tau(U_{B}(n))={\Omega}$, then $u_{B}(n)[0]=0$ and we can take $n^{\prime}=0$, which is in $u_{B}$-normal form. Since $n\geq\min B$, $0\leq n-\min B$. Otherwise, we may apply Lemma 7.9 to $U_{B}(n)$, where in the limit case we set $\iota=0$ and $\Delta=\min B$. ∎
The situation for $i>0$ is more subtle. Recall that we write ${\uparrow}_{i}$ for ${\uparrow}_{\mathcal{D}_{i}}^{\mathcal{D}_{i+1}}$. Then, $\mathcal{D}_{1}$ can compute fundamental sequences up to $1$, but only for numbers of the form ${\uparrow}_{0}n$. Similarly, $\mathcal{D}_{1}$ can compute fundamental sequences up to $2$, but only for numbers of the form ${\uparrow}_{1}{\uparrow}_{0}n$. More generally, $\mathcal{D}_{i}$ can compute fundamental sequences up to $i$, but only to numbers that have been successively upgraded $i$ times.
In order to prove this uniformly, it is convenient to identify a key structural property of $\uparrow$$\mathbb{N}$, i.e., of the image of the upgrade operator. Let $B$ be a base hierarchy and $X\subseteq\mathbb{N}$. We say that $X$ is $B$-digit closed if whenever $n\in X$, $b:={\rm Base}_{B}(n)\in X$, and if $d\in{\rm Digit}_{b}(n)$, it follows that $d\in X$. Obviously, $\mathbb{N}$ is always $B$-digit closed, and the following is readily verified using Lemma 3.5(3).
Lemma 7.11.
Let transformations be from a base hierarchy $B$ to a good successor $C$. Then, if $X\subseteq\mathbb{N}$ is $B$-digit closed, it follows that ${\uparrow}X$ is $C$-digit closed.
The takeaway is that if $B$ computes fundamental sequences up to $i$ for any $n\in X$, where $X$ is $B$-digit closed, then $B^{+}$ computes fundamental sequences up to $i+1$ for any $n\in{\uparrow}X$. The following lemma makes this precise. Below, note that neither $m^{\prime}\in X$ nor $n^{\prime}\in{\uparrow}X$ are required (and, in general, this will not be the case).
Lemma 7.12.
Let transformations be from a base hierarchy $B$ to $C=B^{+}$. Suppose that $X\subseteq\mathbb{N}$ is $B$-digit closed, and $i>0$ and $\Delta\leq\min B$ are such that for every $\in X, mm$ is in $u_{B}$-normal form and there exists $m^{\prime}<m$ in $u_{B}$-normal form such that $u_{B}(m^{\prime})=u_{B}(m)[i-1]$, and such that if $u_{B}(m)$ is a limit, then $m^{\prime}\leq m-\Delta$.
Then, if $n\in X$, there is $n^{\prime}<{\uparrow}n$ such that $n^{\prime}$ is in $u_{C}$-normal form and $u_{C}(n^{\prime})=u_{B}(n)[i]$. Moreover:
  1. (1)
    If $u_{B}(n)$ is a successor, then $n^{\prime}=n-1$.
  2. (2)
    If $u_{B}(n)$ is a limit, then we can take $n^{\prime}\leq n-\Delta$.
Proof.
Proceed by induction on $n$. Let $b={\rm Base}_{B}(n)$, $\xi=u_{C}({\uparrow}n)$, and $c={\rm Base}_{C}({\uparrow}n)$. If $n<\min B$, then we must have $\xi[i]=n-1$, so we can take $n^{\prime}=n-1$. If $n\in B$, then $\xi=\omega$ and we take $n^{\prime}=i=\omega[i]$. Note that the assumption that $u_{B}(\min B)[i-1]=u_{B}(m^{\prime})$ for some $m^{\prime}\leq\min B-\Delta$ yields $i-1\leq\min B-\Delta$, so $i\leq\min B+1-\Delta=\min C-\Delta\leq{\uparrow}n-\Delta$.
Otherwise, let $\zeta=U_{B}(n)$ and note that $\zeta>{\Omega}$, since $n\notin B$. By Proposition 7.8, $\zeta=U_{C}({\uparrow}n)$, and $\xi=\psi(\zeta)$ is in $\psi$-normal form by Lemma 7.5. Then, $\xi[i]=\psi(\zeta^{\prime})$ with $\zeta^{\prime}=\zeta[\iota]$, where either $\tau(\zeta)=\omega$ and $\iota=i$, or else, $\tau(\zeta)={\Omega}$ and $\iota=\xi[i-1]$ (unless $\tau(\zeta)={\Omega}$ and $i=0$, but in this case we can just set $n^{\prime}=0$). Then, consider the following cases.
  1. Case 1
    ($\zeta^{\prime}<{\Omega}$). Note in this case that $\zeta$ is a limit, for otherwise we would have $\zeta=\zeta^{\prime}+1<{\Omega}$, contrary to assumption. Moreover, $\zeta^{\prime}\in\{0,1\}$ by Lemma 6.6(1), and thus $\xi[i]=\psi(\zeta^{\prime})\in\{1,2\}$. Note that since $B$ is a base hierarchy, $\min B\geq 2$, so $\min C\geq\min B+1\geq 3$, and hence we can take $n^{\prime}=\xi[i]<\min C$, which is in $u_{C}$-normal form by default. Since $n\notin B$, $n\geq\min C+1\geq\min B+2$, and hence $n^{\prime}\leq 2\leq n-\min B\leq n-\Delta$.
  2. Case 2
    ($\zeta^{\prime}\geq{\Omega}$). By Lemma 7.11, ${\uparrow}X$ is $C$-closed, given that $X$ is $B$-closed. Thus by the induction hypothesis, $U^{c}_{C}({\uparrow}n)$ satisfies the assumptions of Lemma 7.9(2) regarding $c$-digits. With this in mind, consider two sub-cases according to the cofinality of $\zeta$.
    1. Case 2.1
      ($\tau(\zeta)=\omega)$. In this case, $\zeta^{\prime}=\zeta[i]$. By Lemma 7.9, there is $n^{\prime}$ in $U^{c}_{C}$-normal form with either $n^{\prime}=n-1$ if $\zeta$ is a successor or $n^{\prime}\leq{\uparrow}n-\Delta$ if $\zeta$ is a limit, and such that $U^{c}_{C}(n^{\prime})=\zeta[i]$. If $n^{\prime}<c$, then $n^{\prime}$ is in $u_{C}$-normal form since it is a $c$-digit in $U^{c}_{C}$-normal form. Otherwise, $c\leq n^{\prime}<n$, so ${\rm Base}_{C}(n^{\prime})=c$. Thus, $U_{C}(n^{\prime})=U^{c}_{C}(n^{\prime})=\zeta[i]$, and $n^{\prime}$ is in $U_{C}$-normal form. Moreover, $u_{C}(n^{\prime})=\psi(\zeta[i])$, which is in $\psi$-normal form by Lemma 6.5(2), so Lemma 7.5 implies that $n^{\prime}$ is in $u_{C}$-normal form, as needed.
    2. Case 2.2
      ($\tau(\zeta)={\Omega})$. In this case, $\xi$ must be a limit and $\zeta^{\prime}=\zeta[\xi[i-1]]$. Let $c^{\prime}={\rm Base}_{C}({\uparrow}{n-1})$. We claim that there is $s<c$ such that $s$ is in $u_{C}$-normal form and $u_{C}(s)=\xi[i-1]$. The assumption on $B$ yields $s^{\prime}\leq n-\Delta$ such that $u_{B}(s^{\prime})=u_{B}(n)[i-1]$ and $s^{\prime}$ is in $u_{B}$-normal form. Taking $s={\uparrow}s^{\prime}$, we see by Proposition 7.8 that
      $u_{C}(s)=u_{B}(s^{\prime})=u_{B}(n)[i-1]=u_{C}({\uparrow}n)[i-1]=\xi[i-1]$
      and $s$ is in $u_{C}$-normal form. Moreover, since $s^{\prime}\leq n-\Delta$, we have that $s^{\prime}\leq(n-1)-(\Delta-1)$, so
      $s={\uparrow}s^{\prime}\leq{\uparrow}((n-1)-(\Delta-1))\leq{\uparrow}(n-1)-(\Delta-1)\leq c-\Delta,$
      where the first two inequalities use Lemma 3.1(1) and the third uses Proposition 5.2(2), so $U^{c}_{C}({\uparrow}n)$ satisfies the assumptions of Lemma 7.9 with $\iota:=u_{C}(s)$. Hence, there is $n^{\prime}\leq{\uparrow}n-\Delta$ such that $U^{c}_{C}(n^{\prime})=\zeta[\xi[i-1]]=\zeta^{\prime}$ and $n^{\prime}$ is in $U^{c}_{C}$-normal form. If $n^{\prime}<c$, then it must be in $u_{C}$-normal form. Otherwise, ${\rm Base}_{C}({\uparrow}n)=c\leq n^{\prime}<n$, so ${\rm Base}_{C}(n^{\prime})=c$ and $n^{\prime}$ is in $U_{C}$-normal form. Finally, $u_{C}(n^{\prime})=\psi(\zeta^{\prime})$, which is in $\psi$-normal form by Lemma 6.5. We conclude that $n^{\prime}$ is in $u_{C}$-normal form, as required.∎
Henceforth, we will focus our attention on the greedy dynamical hierarchy $\mathcal{D}=(\mathcal{D}_{i})_{i\in\mathbb{N}}$. Recall that when a dynamical hierarchy is fixed, we may write ${\uparrow}_{i}$ instead of ${\uparrow}_{\mathcal{B}_{i}}^{\mathcal{B}_{i+1}}$. More generally, for $\ell\geq i$, we will define ${\uparrow}_{i}^{\ell}$ by letting ${\uparrow}_{i}^{i}$ be the identity and ${\uparrow}_{i}^{\ell+1}={\uparrow}_{\ell}\circ{\uparrow}_{i}^{\ell}$. As the next proposition shows, as $\ell$ increases, $\mathcal{D}_{\ell}$ can compute more and more of the fundamental sequences of ${\uparrow}_{i}^{\ell}n$.
@@
Let $\mathcal{D}$ be the greedy dynamical base hierarchy. Write $u_{i}$ for $u_{\mathcal{D}_{i}}$ and let $i,j\in\mathbb{N}$. Suppose that $n>0$ is in $u_{i}$-normal form. Then, there is $n^{\prime}<{\uparrow}_{i}^{i+j}n$ such that $n^{\prime}$ is in $u_{i}$-normal form and $u_{i+j}(n^{\prime})=u_{i}(n)[j]$. If moreover $u_{i}(n)$ is a limit, then we can take $n^{\prime}\leq n-i-2$.
Proof.
For $j\in\mathbb{N}$, let $X_{j}={\uparrow}^{i+j}_{i}\mathbb{N}$; we must find suitable $n^{\prime}$ for $n\in X_{j}$. For $j=0$, we have that $X_{0}=\mathbb{N}$, so $n^{\prime}$ can be found by Lemma 7.10, given that $i+2=\min\mathcal{D}_{i}$. Otherwise, assume inductively the claim for all $m\in X_{j}$. Note that $X_{j+1}={\uparrow}_{j}X_{j}$, so Lemma 7.12(2) then yields the desired $n^{\prime}$. ∎
@@
To be precise, we require the following for all $i<I$:
  1. (1)
    if $j<i$, then $t_{j}<t_{i}$;
  2. (2)
    $n_{i}$ is in $u_{t_{i}}$-normal form;
  3. (3)
    $u_{t_{i}}(n_{i})=\xi\llbracket i\rrbracket$;
  4. (4)
    $n_{i}\in{\uparrow}_{t_{i}-i}^{t_{i}}\mathbb{N}$;
  5. (5)
    letting $b_{i}:=t_{i}+2$, we have that
    $t_{i}\geq 2i+[n_{i}]_{b_{i}}+1$, and
  6. (6)
    $n_{i}\leq m_{t_{i}}$.
Conditions (1), (3), and (6) are sufficient to obtain the desired lower bound, but the others are needed to ensure progress of the recursive construction.
We define $n_{i}$ recursively as follows. Set $n_{0}=m_{1}$ and $t_{0}=1$, which are readily checked to satisfy all required properties. Then, suppose that $i+1<I$ and $n_{i}$ has been defined and satisfies (1)-(6). By (4), $n_{i}\in{\uparrow}_{t_{i}-i}^{t_{i}}\mathbb{N}$, and by (5), $t_{i}-i\geq i+1$. Using Proposition 7.13, choose $s_{0}<{\uparrow}_{t_{i}}n_{i}$ in $u_{t_{i}+1}$-normal form such that $u_{t_{i}+1}(s_{0})=u_{t_{i}}(n_{i})[i+1]$, with $s_{0}={\uparrow}_{t_{i}}n_{i}-1$ if $u_{t_{i}}(n_{i})$ is a successor and $s_{0}\leq{\uparrow}_{t_{i}}n-i-3$ if $u_{t_{i}}(n_{i})$ is a limit. Since, by (3), $u_{t_{i}}(n_{i})=\xi\llbracket i\rrbracket$, $u_{t_{i}+1}(s_{0})=\xi\llbracket i+1\rrbracket$.
Now, consider two cases.
  1. Case 1
    ($u_{t_{i}}(n_{i})$ is a successor). Set $n_{i+1}=s_{0}$ and $t_{i+1}=t_{i}+1$. We check that all required properties hold.
    (1) Clearly, $t_{i+1}>t_{i}$.
    (2)-(3) Follow from the fact that $s_{0}$ was chosen to be $u_{t_{i}+1}$-normal form with $u_{t_{i}+1}(s_{0})=\xi\llbracket i+1\rrbracket$.
    (4) Use (4) inductively to write $n_{i}={\uparrow}_{t_{i}-i}^{t_{i}}(x+1)$. Lemma 3.6(3) and backwards induction on $j\leq i$ show that ${\uparrow}_{t_{i}-i}^{t_{i}-j+1}(x+1)={\uparrow}_{t_{i}-i}^{t_{i}-j+1}x+1$, so for $j=0$, we obtain $n_{i+1}+1={\uparrow}_{t_{i}-i}^{t_{i}+1}x+1$, thus $n_{i+1}={\uparrow}_{t_{i}-i}^{t_{i}+1}x$.
    (5) We have that ${\uparrow}_{t_{i}}n_{i}$ is not a multiple of $b_{i+1}$ by Lemma 7.2(1), so $n_{i+1}]_{b_{i+1}}=[{\uparrow}_{t_{i}}n_{i}]_{b_{i+1}}-1=[n_{i}]_{b_{i}}-1$, where the last equality uses Lemma 3.6(1). We thus have
    $t_{i+1}$ $=t_{i}+1\geq 2i+[n_{i}]_{b_{i}}+2=2(i+1)+[n_{i+1}]_{b_{i+1}}+1.$
    (6) From $n_{i}\leq m_{t_{i}}$ and monotonicity (Lemma 3.1(1)), we obtain
    $n_{i+1}={\uparrow}_{t_{i}}n_{i}-1\leq{\uparrow}_{t_{i}}m_{t_{i}}-1=m_{t_{i}+1}=m_{t_{i+1}}.$
  2. Case 2
    ($u_{t_{i}}(n_{i})$ is a limit). Then, $s_{0}\leq n_{i}-i-3$. For $j>1$, define $s_{j}={\uparrow}_{t_{i}+1}^{t_{i}+j}s_{0}$. Set $t_{i+1}=t_{i}+i+3$ and $n_{i+1}=s_{i+2}$.
    (1) Clearly, $t_{i+1}>t_{i}$.
    (2) This holds in view of Proposition 7.8, given that $s_{0}$ is in $u_{t_{i}+1}$-normal form, so induction shows that each $s_{j}$ (including $s_{i+2}=n_{i+1}$) is in $u_{t_{i}+j}$-normal form.
    (3) Proposition 7.8 and induction on $j$ yield
    $\xi\llbracket i+1\rrbracket=u_{t_{i}}(n_{i})[i+1]=u_{t_{i}+1}(s_{0})=u_{t_{i}+j}(s_{j})=u_{t_{i+1}}(n_{i+1}).$
    (4) This holds since $n_{i+1}={\uparrow}_{t_{i}+2}^{t_{i}+i+3}s_{1}$. [^4] Note that we purposefully use $s_{1}$ and not $s_{0}$ in this step, since the latter would force us to set $t_{i+1}=t_{i}+i+2$, which is too small for the subsequent (5).
    (5) By Lemma 6.6(2), $u_{t_{i}}(n_{i})[i+1]=\omega{\alpha}+r$ for some $r\leq i+1$. Lemma 7.12(1) applied $r$ times yields $n_{i+1}=p+r$ with $u_{t_{i+1}}(p)=\omega{\alpha}$, and thus by Lemma 7.2(1), $p$ is a multiple of $\min\mathcal{D}_{t_{i+1}}=b_{i+1}$. We conclude that $n_{i+1}]_{b_{i+1}}=r\leq i+1$. Then, since $t_{i}\geq 2i+1$,
    $t_{i+1}=t_{i}+i+3\geq 3i+4=3(i+1)+1\geq 2(i+1)+[n_{i+1}]_{b_{i+1}}+1.$
    (6) We show inductively using Lemma 3.1(1) that for $0\leq j\leq i+2$, $s_{j}\leq m_{t_{i}+j}-i-2+j$. For the base case, we use (6) to see that $n_{i}\leq m_{t_{i}}$, and thus using Lemma 3.1(1),
    $s_{0}\leq{\uparrow}_{t_{i}}n_{i}-i-3\leq{\uparrow}_{t_{i}}m_{t_{i}}-i-3=m_{t_{i}+1}-i-2.$
    For the inductive step,
    $s_{j+1}$ $={\uparrow}_{t_{i}+j}s_{j}\leq{\uparrow}_{t_{i}+j}(m_{t_{i}+j}-i-2+j)$
    $\leq{\uparrow}_{t_{i}+j}m_{t_{i}+j}-i-2+j=m_{t_{i}+j+1}-i-2+j+1.$
    Setting $j=i+2$, we see that $n_{i+1}=s_{i+2}\leq m_{t_{i}+i+2}=m_{t_{i+1}}$.∎
8. Finite Base Hierarchies
The proof of Theorem 2 uses a dynamical hierarchy whose base hierarchies (aside from the first) are infinite. One may wonder if the use of infinite base hierarchies is essential. The answer is no; here we will briefly sketch how our constructions should be modified to show this. Below, say that a finite dynamical hierarchy is a dynamical hierarchy $\mathcal{B}$ such that $\mathcal{B}_{i}$ is finite for all $i$.
Theorem 2.